OpenStudy (anonymous):

Evaluate the expression for a = –1 and b = 5. 5a – 7

OpenStudy (hoblos):

replace a by (-1) and b by (5) 5a – 7 = 5(-1) -7 = -5-7 = -12 but where is the b in the b in the expression!!!